40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15017     Entry time: Wed Nov 6 19:26:57 2019
Author: gautam 
Type: Update 
Category: PSL 
Subject: Some PSL cable admin 

Koji and I taked about cleaning up some of the flaky cable situation on the PSL table a while ago. The changes were implemented and are documented in Attachment #1. Now the Pomona box between the Thorlabs HV Driver and the NPRO head is sitting on the PSL table (sandwiched between some teflon pieces I found in cabinet S4 along the south arm), and the cables between these two devices are better strain relieved. I turned off the Thorlabs HV supply while working on the PMC table. The IMC could be locked after this work. Probably won't solve the long standing FSS mysteries but probably can't hurt.

Unrelated to this work: I also removed a Bias tee that was just hanging out on top of the FSS electronics, which was used for the modeSpec project.

Attachment 1: PSLcableAdmin.jpg  476 kB  | Hide | Hide all
PSLcableAdmin.jpg
ELOG V3.1.3-